# HoneySat: A Network-based Satellite Honeypot Framework

### v1

@@ 5.1 Satellite Simulator Implementation
We implemented 11 sensors, 4 subsystems, 1 satellite state, and 1 interface, based on our 6 simulations. These modules are generic enough so that we do not have to heavily modified them to support other satellites. Overall, satellite architectures, subsystems, and sensors are varied, so modularity is necessary when constructing a simulation framework. Many of the basic simulations and corresponding sensors can be shared, especially in a honeypot where attackers lack detailed information that would enable them to distinguish between small differences in sensor values. Further details on the simulations can be found in A.2
5.2 Generic CSP Mission Honeypot
Following our design we implemented HoneySat to support the CSP ecosystem. This implementation provides a generic CSP-based mission honeypot that can be customized to simulate a specific CPS-based satellite missions.
5.2.1 Ground Segment Implementation
We now describe how we implemented the five ground segment simulations designed in Sec. 4.3. 1) the Exposed Entry Points, 2) the Ground Software, 3) the Radio Proxy, 4) the Ground Configuration, and 5) Logging Repository.
1) Exposed Network Protocols. We implemented four exposed entry points. A VNC server, a Telnet server, a Web server, and ZeroMQ-based access to the CSP Network. We implemented the VNC server using TigerVNC [51] and used PyZMQ [25] for the Raw Packet Access. The Telnet server was implemented using Python to expose the command-line interface of GSS on the network. The web interface allows for the customization of text presented to the attacker through configuration. Using the above entry points, adversaries can interact with the GSS and the space segment simulations. Among these, the web interface serves as a measure to attract attackers and serves as an initial entry point for attackers. We developed a simple yet convincing web application, presented as an early version of a web interface currently under development. After the login, the attacker would be welcomed by a dashboard that displays information about the satellite and the latest output of the Ground software command line interface. Additionally, the web application provides access to documentation and the VNC server (Fig. 4), facilitating the attacker in finding the other entry points.
@@ 5.2.1 Ground Segment Implementation
2) Ground Software. To implement the MCS software for our CSP-based honeypot, we require an open-source MCS for the CSP ecosystem. We leveraged the SUCHAI mission control software (SUCHAI MCS). This software is a node on the CSP network that has a command line-based interface that allows basic MCS functionality like scheduling telecommands or saving downlinked telemetry. This is a good choice for the generic CSP MCS as it can be passed off as the CLI of commercial ground segment appliances like the GomSpace NanoCom MS100 [23]. In this generic honeypot, we can set mission-specific parameters through the ground configuration. Additionally, we provide a desktop environment accessible via VNC, complete with supporting applications such as configured Gpredict for the satellite mission. For certain deployments, this environment also includes a mission-specific GUI tool, both of which appear as open windows on the desktop.
3) Radio Simulator. The Radio Simulator was implemented as a ZMQ Hub. The ZMQ Hub functions as a router for the CSP network. Simulating both the ground part of the CSP network including the ground station and space segment. Depending on whether the satellite is currently passing, packets to the space segment are forwarded to the Flight Software Runtime. We decided to use ZeroMQ as it is the protocol to use for transmitting CSP over TCP/IP [64].
4) Ground Configuration. We implemented the ground configuration on each component, and the settings can be modified through a Docker compose file explained in Appendix A.3.
@@
5.2.2 Space Segment Implementation
1) Flight Software Runtime Although most of the flight software frameworks presented in Section 2.4 are open-source projects, it is difficult to find a real satellite flight software implementation that is available for study and modification [81]. For this reason, we use the University of Chile’s SUCHAI Flight Software (SUCHAI FS) [24], which has been deployed in four satellite missions [21]. The SUCHAI FS provides Flight software based on the CSP ecosystem that already features a mode where it can be run on a POSIX OS.
2) Flight Software Services. The SUCHAI FS uses libCSP [63] and thus features the libCSP default services that present a generic attack surface. Furthermore, there is a set of commands commonly used in satellites as a basic set of tasks to control common satellite operations such as communications. There are 96 of these commands and parameters that are parsed from CSP packets using a text-based protocol.
@@ 5.2.2 Space Segment Implementation
4) Logging Repository. Like the ground segment’s logging repository (Sec 5.2.1), The Satellite Simulator and the FS are connected to the MongoDB instance using a MongoDB client.
Each logged data entry is created with customized tagging and a timestamp. The collected data is structured JSON [32] format and stored in the MongoDB database.
Finally, we containerized and hardened this version of the honeypot implementation for secure and easy deployment, which we describe in Appendix A.3.
6 Evaluation
Having described HoneySat’s implementation in Sec. 5, we now evaluate it. This section lists three experimental questions designed to test HoneySat’s alignment with our design objectives. Next, we present four sets of experiments that provide empirical evidence confirming that our design objectives have been achieved within HoneySat. We describe each experiment’s environment, methodologies, and results.
6.1 Experimental Questions
These questions aim to determine whether or not HoneySat’s meets the design objectives outlined in Sec. 4.1.
@@ 1. Q-1
    Since capturing data on varied techniques is the purpose of any honeypot, we explore the capabilities of HoneySat, as described in Sec. 4. This question is related to design objective DO-1 and is addressed in Sec. 6.2 and 6.4.
@@ 2. Q-2
    After enticing an adversary HoneySat must keep its true nature hidden, we need to simulate the satellite’s communication and physics characteristics described in Sec. 2.4. This question relates to design objective DO-2 and is answered in Sec. 6.3.
@@ 3. Q-3
    Can HoneySat simulate different SmallSat missions?
    HoneySat’s customization is important because it would allow users of our framework to implement their own honeypots. This question relates to design objective DO-3, and we answer it in Sec. 6.5.
To answer the above research questions we conducted four experiments. First, in Sec. 6.2 we craft multiple attacks in a controlled environment to quantify the level of interaction that HoneySat provides. Second, in Sec. 6.3 we conduct a survey with experienced satellite operators to evaluate HoneySat’s realism. Third, in Sec. 6.4 we deploy HoneySat and expose it to the Internet to test its deception capabilities. Fourth and final, in Sec. 6.5 we add an additional protocol ecosystem to HoneySat to test its extensibility potential.
6.2 TTP Interaction Experiment
This experiment seeks to answer Q-1 by quantifying the different interactions that HoneySat provides. To achieve this, we leveraged the SPACE-SHIELD matrix (Version 2.0) [4] discussed in Sec. 2.6. SPACE-SHIELD provides a collection of adversary tactics and techniques for space systems. In this experiment, we determined the number of tactics and techniques that HoneySat supports. SPACE-SHIELD consists of 14 tactics and 62 techniques. However, not all of them are applicable to a virtual, network-based honeypot such as HoneySat. For example, the technique Compromise Hardware Supply Chain involves “replacing a hardware component in the supply chain with a custom or counterfeit part” which is out of the scope of a virtual honeypot. Taking this into consideration, 14 tactics, and 33 techniques are applicable to HoneySat.
@@ 6.2 TTP Interaction Experiment
Experiment Methodology. We designed one interaction or exploit for each of the applicable 33 techniques for our honeypot. The interactions involved ground segment simulations such as the web interface. The exploits were simple, using one TC, or complex with multiple TCs involved.
Experiment Results. We crafted 16 exploits and 17 interactions following the above methodology. For example, the technique T1007 - System Service Discovery’s description reads “Adversaries may try to gather information about registered local system services. Adversaries may obtain information about services using tools and OS utility commands. Adversaries may use the information from System Service Discovery during automated discovery to shape follow-on behaviors, including whether or not the adversary fully infects the target and/or attempts specific actions.”
@@
Similarly, technique T2014 - Backdoor Installation reads “An attacker can interfere with the hardware or the software, integrating or modifying the existing software, hardware configuration, or the transponder configuration to permit himself future access to the resource”. Therefore, we designed the following exploit to upload a malicious script to the satellite software:
1tm_send_file backdoor.sh 1
21: obc_system ./recv_files/backdoor.sh
Due to space limitations, we do not describe all the exploits and interactions here. However, the complete exploit and interaction list is available in Table 6 in Appendix A. The overall results for this experiment are depicted in Table 3. The key findings of our experiment are shown below, providing strong evidence for answering question Q-1 in the affirmative and design objective DO-1 as achieved.
Key findings Q-1 HoneySat supports 86.8% of the SPACE-SHIELD matrix techniques possible in a virtual satellite honeypot. HoneySat supports 100% of the SPACE-SHIELD matrix tactics.
@@
Table 3: Tactics and techniques supported by HoneySat compared to the ones included in the SPACE-SHIELD matrix that are applicable to virtual honeypots.
6.3 SmallSat Operators Survey
Evaluating the realism of a satellite honeypot is challenging for two main reasons. First, as we discussed in Sec. 1, satellites, including SmallSats are very diverse. Second, there is no established metric or tool, such as Nmap’s OS detection [49], to quantify the level of realism of our honeypot.
To evaluate the realism and deception capabilities of HoneySat, we surveyed experienced, real-world SmallSat operators. Because operators interact with real-world SmallSat missions on a daily basis they are experts and thus are the best population to rigorously evaluate HoneySat.
@@ 6.3 SmallSat Operators Survey
Telecommand Scheduling Mission Control SW 6
Pass Prediction Ground Station Control SW 6
Antenna Positioning Ground Station Control SW 6
EPS Telemetry Satellite Simulator 6
Temperature Telemetry Satellite Simulator 6
ADCS Telemetry Satellite Simulator 6
RGB Camera Telecommand Satellite Simulator 6
Ping Test Radio Simulator 6
Table 4: Summary of questions in Section 6.3, satellite honeypot operation tasks of the survey (48 total).
6.3.1 Survey Structure
We divided our survey into five sections. The first section collected essential background information about the participants, while the second part focused on their professional experience. In this second section, participants were asked whether they worked in industry, government, or academia and to self-report their cybersecurity skills as well as satellite-related skills (e.g., satellite hardware engineering). The third section of the survey probed participants’ satellite operation experience, including how many missions they had operated and which tools they used.
The largest section of the survey included 48 questions about satellite honeypot operation tasks. Participants watched eight 1–2 minute recordings of HoneySat ’s VNC view (Fig. 4), each showcasing the ground and satellite personality of a real CSP-based CubeSat mission (based on the SUCHAI-2 CubeSat [34]). Every recording highlighted a different feature of HoneySat by replicating a real-world satellite mission operation, as discussed in Section 2. After each recording, participants answered questions designed to assess how realistic they found that particular aspect of the honeypot. Table 4 lists all of the satellite operations demonstrated in these recordings. Finally, once participants had seen multiple elements of our honeypot, the survey concluded with an overall evaluation section. Both the honeypot operation tasks and the overall evaluation used 5-point Likert scale [55] questions to measure participants’ positive or negative reactions.
6.3.2 Participants
We conducted the survey via Qualtrics and distributed the survey directly to active SmallSat mission operators from previously identified missions. In total, we received responses from 14 satellite operators who have between 1 and 5 years of experience operating satellites and have operated between 2 and 7 unique missions. In terms of demographics, 21.4% (3/14) of participants were female, 71.4% (10/14) male and 7.1% (1/14) preferred not to answer. 57.1% (8/14) belonged to the 18-24 age group, 35.7% (5/14) to 25-34 and 7.1% (1/14) to 35-44. 28.5% (4/14) of participants’ highest level of education was high school, 35.7% (5/14) bachelor’s degree, 14.2% (2/14) master’s degree and 21.4% (3/14) Ph.D. In regards to geographic location, 78.5% (11/14) of participants were located in Europe, 14.2% (2/14) in North America and 7.1% (1/14) in South America.
@@ 6.3.3 Methodology and Key Results
In the survey, we aimed to evaluate three key aspects of our honeypot. First, whether the ground segment simulations are realistic; second, whether the space segment simulations are realistic; and third, whether HoneySat, as a whole, provides a convincing and realistic system.
Before describing the results it is important to emphasize that the participants were informed that the recordings showed a simulation of a satellite mission and not of a real mission.
Ground Segment Realism. To understand if HoneySat’s ground segment is realistic we showed participants recordings of different satellite operations (discussed in Sec. 2.3) performed with HoneySat, to showcase the ground segment components listed in Table 4.
@@
For example, one of the recordings shows the pass prediction operation which involves calculating when and where a satellite will be visible or within range of a specific ground station. This operation is performed using the ground station control software. After the participants watched a recording of this operation, we asked them to rate their perceived level of realism. 85.7% of participants agreed or strongly agreed that the pass prediction operation done in HoneySat resembles that of a real mission.
@@
Another relevant operation is the telecommand scheduling operation which involves the planning and organization of the commands to be sent to the satellite during a pass. 42.8% of participants strongly agreed that the telecommand scheduling operation performed in HoneySat resembles that of a real mission mentioning the use of a terminal-based mission control software. Interestingly, 35.7% of participants neither agreed nor disagreed, citing that they do not use a command line tool but instead a GUI. These results indicate that some of the surveyed operators use a command line-based mission control software while others use a GUI-based confirming the diversity of means of operation in Sec. 2.3.
In summary, according to our survey, most participants perceive the ground segment simulated by HoneySat as comparable to that of a real satellite mission. While a few participants mentioned that they expected to see a GUI-based mission control software instead of a command line-based one, this is something that we expected, as the means of satellite operations vary greatly among missions. Nevertheless, thanks to HoneySat’s modularity HoneySat can be extended to use a GUI-based mission control software such as YAMCS [1].
Space Segment Realism. In a similar manner, to understand if HoneySat’s space segment is realistic, we showed participants recordings of different satellite operations that make use of HoneySat’s space segment simulations (discussed in Sec. 4.4) to showcase multiple components of the space segment listed in Table 4. For example, one of the recordings shows the EPS Telemetry operation which involves the collection of data from the Electrical Power Subsystem discussed in Sec. 2.4. This operation is performed using the mission control software to send telecommands and download the latest EPS telemetry. After the participants watched a recording of this operation performed by HoneySat, we asked them to rate their perceived level of realism of the telemetry output shown during the operation. 57.1% of participants agreed or strongly agreed that the telemetry shown in the EPS Telemetry operation resembles that of a real mission mentioning realistic EPS values such as voltage, and temperature. 14.2% neither agreed nor disagreed mentioning that the EPS temeletry was presented in a different format. Finally, 28.5% disagreed or strongly disagreed mentioning that the telemetry included less values than the mission that they had operated. These results indicate that the EPS telemetry generate by the Satellite Simulator is considered realistic by the majority of participants.
@@
We also asked participants to asses the ping test operation. This operation is a diagnostic procedure used to verify the communication link between a ground station and the satellite. It involves sending CSP packets between two CSP nodes, namely the ground segment and the spacecraft. After the participants watched a recording of this operation performed on HoneySat, we asked them to rate their perceived level of realism. 64.2% of participants agreed or strongly agreed that the ping test operation resembles that of a real mission citing the realistic response times. 21.4% of participants neither agreed nor disagreed stating that in their experience the ping test also downloads additional telemetry. Finally, 14.2% disagreed or strongly disagreed noting that in their mission the ping command has a different name. The results of the ping test operation indicate that HoneySat’s radio simulator (which controls the communication between the ground and space segments) is considered to realistically simulate the communication of a real satellite mission by the majority of participants. A significant key finding of our survey is shown below. Overall, the results obtained provide convincing evidence for answering question Q-2 in the affirmative and design objective DO-2 as achieved.
@@
Key finding Q-2 71.4% of surveyed satellite operators agreed or strongly agreed with the following statement: “If I did not know this was a honeypot simulation of a CubeSat satellite mission, I would have believed it to be an actual CubeSat satellite mission.”
6.4 Internet Interaction Experiment
This experiment explores HoneySat’s capabilities to entice external actors in the wild by deploying our honeypot over the Internet. To accomplish this, we leveraged the Shodan search engine. Shodan is a search engine for Internet-connected devices that scans the whole Internet and indexes exposed servers and their TCP ports. Shodan then reads the banner information for each port. For example, it gathers the web headers and Telnet login banners [40]. Additionally, Shodan tags servers as honeypots if the servers have too many open ports [37].
Experiment Description. We deployed and configured five HoneySat instances over the Internet and exposed TCP port 24 for the Telnet server and 80 for the web interface so that anybody can interact with it and allow Shodan to index our server and its banners. We deployed four HoneySat instances in total.
Honeypot
Source

### v2

@@ V-A Satellite Simulator Implementation
We implemented 11 sensors, 4 subsystems, 1 satellite state, and 1 interface, based on our 6 simulations. These modules are generic enough so that we do not have to heavily modified them to support other satellites. Overall, satellite architectures, subsystems, and sensors are varied, so modularity is necessary when constructing a simulation framework. Many of the basic simulations and corresponding sensors can be shared, especially in a honeypot where attackers lack detailed information that would enable them to distinguish between small differences in sensor values. Further details
can be found in Appendix A-B.
V-B Generic CSP Mission Honeypot
Following our design, we implemented HoneySat to support the CSP ecosystem. This implementation provides a generic CSP-based mission honeypot that can be customized to simulate a specific CPS-based satellite mission.
V-B1 Ground Segment Implementation
We now describe how we implemented the five ground segment simulations designed in Sec. IV-C. 1) the Exposed Entry Points, 2) the Ground Software, 3) the Radio Proxy, 4) the Ground Configuration, and 5) Logging Repository.
1) Exposed Network Protocols.
We implemented four exposed entry points. A VNC server, a Telnet server, a Web server, and ZeroMQ-based access to the CSP Network. We implemented the VNC server using TigerVNC [64] and used PyZMQ [65] for the Raw Packet Access. The Telnet server was implemented using Python to expose the command-line interface of the Ground Segment Software (GSS) on the network. The web interface allows for the customization of text presented to the attacker through configuration. Using the above entry points, adversaries can interact with the GSS and the space segment simulations.
@@ V-B1 Ground Segment Implementation
2) Ground Software. To implement the MCS software for our CSP-based honeypot, we require an MCS that supports the CSP ecosystem. We leveraged the SUCHAI mission control software (SUCHAI MCS). This software is a node on the CSP network that has a command line-based interface that allows basic MCS functionality like scheduling telecommands or saving downlinked telemetry. This is a good choice for the generic CSP MCS as it can be passed off as the CLI of commercial ground segment appliances like the GomSpace NanoCom MS100 [46]. Given HoneySat being a generic honeypot, we can set mission-specific parameters through the ground configuration. Additionally, we provide a desktop environment accessible via VNC, complete with supporting applications such as a configured Gpredict for the satellite mission. For certain deployments, this environment also includes a mission-specific GUI tool, both of which appear as open windows on the desktop.
3) Radio Simulator. The Radio Simulator was implemented as a ZMQ Hub [66]. The ZMQ Hub functions as a router for the CSP network simulating both the ground part of the CSP network including the ground station and the space segment. Depending on whether the satellite is currently passing, packets to the space segment are forwarded to the Flight Software Runtime.
We decided to use ZeroMQ as it is the protocol to use for transmitting CSP packets over TCP/IP [66].
4) Ground Configuration. We implemented the ground configuration as settings that can be modified through a Docker compose file explained in Appendix A-C.
@@
V-B2 Space Segment Implementation
We now describe how we implemented the four space segment simulations designed in Sec. IV-D. 1) the Flight Software Runtime, 2) the Flight Software Services, 3) the Satellite Personality,and 4) the Logging Repository.
1) Flight Software Runtime. Although most of the flight software frameworks presented in Section II-D are open-source projects, it is difficult to find a real satellite flight software implementation that is available for study and modification [12]. Thus, we use the University of Chile’s SUCHAI Flight Software (SUCHAI FS) [52], which has been deployed in four satellite missions [68]. The SUCHAI FS provides flight software based on the CSP ecosystem that already features a mode where it can be run on a POSIX OS.
2) Flight Software Services. The SUCHAI FS uses libCSP [55] and thus features the libCSP default services that present a generic attack surface. Furthermore, there is a set of commands commonly used in satellites as a basic set of tasks to control common satellite operations such as communications. There are 96 of these commands and parameters that are parsed from CSP packets using a text-based protocol. We modified the command handlers to use the Satellite Simulator instead of commanding subsystems and querying real sensors.
@@ V-B2 Space Segment Implementation
4) Logging Repository. Like the ground segment’s logging repository (Sec V-B1), the Satellite Simulator and the FS are connected to the MongoDB instance using a MongoDB client.
Each logged data entry is created with a customized tag and a timestamp. The collected data is structured JSON [69] format and stored in the MongoDB database.
Finally, we containerized and hardened this version of the honeypot implementation for secure and easy deployment, which we describe in Appendix A-C.
VI Evaluation
In order to evaluate HoneySat we start by listing three experimental questions designed to test its alignment with our design objectives. Next, we present four sets of experiments that provide empirical evidence confirming that our design objectives have been addressed as intended. We then describe each experiment’s environment, methodologies, and results.
VI-A Experimental Questions
The following questions aim to determine if HoneySat meets the design objectives outlined in Sec. IV-A.
@@ 1. Q-1
    Since capturing data on varied techniques is the purpose of any honeypot, we explore the capabilities of HoneySat, as described in Sec. IV. This question is related to design objective DO-1 and is addressed in Sec. VI-B.
@@ 2. Q-2
    After enticing an adversary, HoneySat must keep its true nature hidden. Thus, HoneySat needs to simulate the satellite’s communication and physics characteristics described in Sec. II-D. This question relates to design objective DO-2 and is answered in Sec. VI-C and VI-D.
@@ 3. Q-3
    Can HoneySat be extended to simulate different SmallSat missions and support hardware-in-the-loop simulations?
    HoneySat’s customization is important because it would allow users of our framework to implement their own honeypots and even support satellite hardware integration. This question relates to design objective DO-3, and we answer it in Sec. VI-E and Sec. VI-F.
To answer the above research questions, we conducted five experiments. First, in Sec. VI-B we craft multiple attacks in a controlled environment to quantify the level of interaction that HoneySat provides. Second, in Sec. VI-C we conduct a survey with experienced satellite operators to evaluate HoneySat’s realism. Third, in Sec. VI-D we deploy HoneySat and expose it to the Internet to test its deception capabilities. Fourth, in Sec. VI-E we test HoneySat’s extensibility by integrating a completely different flight software. Fifth and final, in Sec. VI-F we test HoneySat’s support for hardware-in-the-loop operations.
VI-B TTP Interaction Experiment
This experiment seeks to answer Q-1 by quantifying the interactions provided by HoneySat. To achieve this, we leveraged the SPACE-SHIELD matrix (Version 2.0) [63] discussed in Sec. II-F. SPACE-SHIELD provides a collection of adversary tactics and techniques for space systems. In this experiment, we determined the number of tactics and techniques that HoneySat supports. SPACE-SHIELD consists of 14 tactics and 62 techniques. However, not all of them are applicable to a virtual, network-based honeypot such as HoneySat. For example, the technique Compromise Hardware Supply Chain involves “replacing a hardware component in the supply chain with a custom or counterfeit part” which is out of the scope of a virtual honeypot. Taking this into consideration, 14 tactics are applicable to HoneySat. From these tactics, HoneySat can feasibly offer up to 38 techniques as possible interactions for adversaries.
@@ VI-B TTP Interaction Experiment
Experiment Methodology. We designed one interaction or exploit for each of the 38 applicable techniques for our honeypot. The interactions involved ground segment simulations such as the web interface. The exploits were simple, using one TC, or complex with multiple TCs involved.
For example, the technique System Service Discovery (ID T1007)’s description reads “Adversaries may try to gather information about registered local system services. Adversaries may obtain information about services using tools and OS utility commands. Adversaries may use the information from System Service Discovery during automated discovery to shape follow-on behaviors, including whether or not the adversary fully infects the target and/or attempts specific actions.”
@@
Experiment Results. We successfully crafted and ran an interaction or exploit on HoneySat for 33 techniques out of the feasible 38 as depicted in Table III. The remaining 5 techniques were not implemented due to limitations in HoneySat’s implementation. For example, the Retrieve TT&C master/session keys (T2015.002) technique depends on a cryptographic protocol implementation which HoneySat’s flight software currently does not support. This is not surprising, as satellite flight software rarely if ever does [12].
Due to space limitations, we do not describe all the exploits and interactions here. However, the complete exploit and interaction list is available in Table VIII in Appendix A.
The key findings of our experiment are shown below, providing strong evidence for answering question Q-1 in the affirmative and design objective DO-1 as achieved.
Key findings Q-1 HoneySat supports 87% of the SPACE-SHIELD matrix techniques possible in a virtual satellite honeypot. HoneySat supports 100% of the SPACE-SHIELD matrix tactics.
TABLE III: Tactics and techniques supported by HoneySat compared to the ones included in the SPACE-SHIELD matrix that are applicable to virtual honeypots.
@@
VI-C SmallSat Operators Survey
Evaluating the realism of a satellite honeypot is challenging for two main reasons. First, as discussed in Sec. I, satellites, including SmallSats, are very diverse. Second, there is no established metric or tool, such as Nmap’s OS detection [70], to quantify the level of realism of our honeypot.
To evaluate the realism and deception capabilities of HoneySat, we surveyed experienced SmallSat operators. Because operators interact with real-world SmallSat missions on a daily basis they are experts and thus are the best population to rigorously evaluate HoneySat.
TABLE IV: Summary of questions in Section VI-C, satellite honeypot operation tasks of the survey.
@@ VI-C SmallSat Operators Survey
Telecommand Scheduling Mission Control SW 5
Pass Prediction Ground Station Control SW 5
Telemetry Download Satellite Simulator 7
Ping Test Radio Simulator 5
VI-C1 Survey Structure
We divided our survey into five sections. The first section collected the participants’ background information including demographic data. The second section focused on their professional experience, including whether they worked in industry, government, or academia and to self-report their satellite-related and cybersecurity skills (e.g., satellite hardware engineering). The third section probed participants’ satellite operation experience, including how many missions they had operated and which tools they used.
The fourth survey section involved giving participants hands-on access to a live instance of HoneySat. In this section, participants were asked to perform 4 hands-on tasks that correspond to specific satellite mission operations using a live instance of HoneySat as depicted in Fig. 4. Each hands-on task was designed to feature different components of HoneySat by replicating a real-world satellite mission operation, as discussed in Section II. After participants completed each hands-on task, they answered questions designed to assess how realistic they found different aspects of the honeypot. Table IV lists the satellite operations included in the survey.
Finally, in section five, once participants had interacted with multiple elements of our honeypot via these tasks, the survey concluded with participants answering questions about HoneySat’s overall realism and deception. Questions related to both the honeypot operation tasks and the overall evaluation used a 5-point Likert scale [71] to measure the participants’ reactions.
VI-C2 Participants
We conducted the survey via Qualtrics and Zoom videoconferencing and distributed the survey directly to experienced SmallSat mission operators from previously identified missions. In total, we received responses from 10 satellite operators who have between 1 to 10 years of experience operating satellites and have operated between 1 to 5 unique missions. In terms of demographics, 20% (2/10) of the participants were female and 80% (8/10) male. 30% (3/10) belonged to the 18-24 age group, 40% (4/10) to the 25-34 age group, and 30% (3/10) to the 35-44 age group. In regards to geographic location, 40% (4/10) of participants were located in Europe, 20% (2/10) in North America and 40% (4/10) in South America.
@@ VI-C3 Methodology and Key Results
In the survey, we evaluated three key aspects of our honeypot. First, whether the ground segment simulations are realistic; second, whether the space segment simulations are realistic; and third, whether HoneySat, as a whole, provides a convincing and realistic SmallSat mission simulation.
Before describing the results, it is important to emphasize that the participants were informed that the system they interacted with was a simulation of a satellite mission and not a real one. This is because it was not feasible to synchronize participants’ availability to take the survey with the timing of the real satellite’s pass, which may happen only a few times per day and lasts only a few minutes.
@@
Ground Segment Realism. To understand if HoneySat’s ground segment is realistic, participants interacted with HoneySat by performing different satellite operations (discussed in Sec. II-C), to showcase the ground segment components listed in Table IV. One of these tasks is the pass prediction operation which involves calculating when and where a satellite will be within the communication range of a specific ground station. This operation is performed using the ground station control software. After the participants performed this operation on HoneySat, we asked them to rate their perceived level of realism. 90% of the participants strongly agreed and 10% agreed that the pass prediction operation they performed on HoneySat resembles that of a real mission.
@@
Another relevant operation is the telecommand scheduling operation which involves the planning and queuing of the commands to be sent to the satellite during a pass. Again, after performing this operation on HoneySat, 90% of the participants strongly agreed and 10% agreed that the telecommand scheduling operation performed resembles that of a real mission mentioning the use of a command line interface-based mission control software as a contributing factor.
In summary, according to these results, the vast majority of participants, who are experienced satellite operators, perceive the ground segment simulated by HoneySat as highly resembling a real satellite mission.
Space Segment Realism. Similarly, to understand HoneySat’s space segment realism level, we gave participants hands-on access to a live instance of HoneySat and asked them to perform different operations that make use of HoneySat’s space segment simulations (discussed in Sec. IV-D).
For example in the telemetry download operation, participants issued multiple telecommands to HoneySat’s space segment simulation to download a plethora of telemetry data, including EPS, temperature and ADCS data, which was generated in real time by our Satellite Simulator.
After the participants performed the telemetry download operation, we asked them to rate their perceived level of realism of the telemetry output shown during the operation. 90% of the participants agreed or strongly agreed that the telemetry shown during the live telemetry download operation resembles that of a real mission, with the temperature telemetry achieving a 100% strongly agree response rate with some participants mentioning that the temperature values resemble a normal operation and align with the expected temperature values of a satellite. For the EPS telemetry, 80% strongly agreed, 10% agreed, and the last 10% neither agreed nor disagreed when asked if they perceived the EPS telemetry as realistic. In this case, one participant pointed out that in their specific mission they have more batteries, which is expected as satellite missions are highly diverse as we discussed in Sec. II-C. These results indicate that the telemetry generated by the Satellite Simulator is considered highly realistic by the vast majority of participants.
@@
We also asked the participants to assess the ping test operation. This operation is a diagnostic procedure used to verify the communication link between a ground station and the satellite. It involves sending CSP packets between two CSP nodes, namely the ground segment and the spacecraft. After the participants completed this operation’s hands-on task on HoneySat, we asked them to rate their perceived level of realism. 70% of the participants strongly agreed and 30% agreed that the ping test operation resembles that of a real mission with some participants citing that communication was only possible when the satellite was performing a pass.
The results of the ping test operation indicate that HoneySat’s radio simulator (which controls the communication between the ground and space segments) is considered to realistically simulate the communication of a real satellite mission by all participants.
In summary, according to these results, the vast majority of participants, who are experienced satellite operators, perceive the space segment simulated by HoneySat as highly resembling a real satellite mission.
HoneySat’s Overall Realism. Finally, to understand HoneySat’s overall realism, we asked the participants a series of overall evaluation questions after they had interacted with HoneySat. For example, we asked participants to evaluate the realism of the telecommands used in all hands-on tasks. 70% strongly agreed, 20% agreed and 10% neither agreed nor disagreed that HoneySat’s telecommands are realistic. Additionally, when asked if they would be able to distinguish between HoneySat’s satellite mission simulation and a real mission, 70% strongly agreed, 20% agreed and 10% neither agreed nor disagreed that they would not be able to and again to emphasize that these are experienced satellite operators.
Overall, our survey results provide convincing evidence for answering question Q-2 in the affirmative and design objective DO-2 as achieved.
@@
Key finding Q-2 90% of surveyed experienced satellite operators agreed or strongly agreed with the following statement: “I would not be able to distinguish between HoneySat’s satellite honeypot system from the real CubeSat satellite mission it is based on.”
VI-D Internet Interaction Experiment
This experiment explores HoneySat’s capabilities to entice external actors to interact with it by deploying our honeypot over the Internet.
Experiment Methodology. We leveraged HoneySat’s customization and deployability features to deploy five instances of HoneySat over the Internet and exposed the necessary TCP ports used by HoneySat’s exposed network protocols discussed in Sec. IV-C, e.g., Telnet’s port 23.
Initially, we deployed our first HoneySat instance to simulate the SUCHAI-2 SmallSat on the premises of the University of Chile, where the mission’s real ground segment is. This on-premises deployment provides the most convincing environment to adversaries. After coordinating with the University of Chile’s local stakeholders for three months, we deployed HoneySat on an onsite server using a University of Chile IP address starting on July 2024.
Although we went to great lengths to make our deployment as realistic as possible, after a few months, our honeypot did not entice any external actors. Therefore, in January 2025, we deployed four additional HoneySat instances on cloud servers. We configured two instances to simulate NASA’s Advanced Composite Solar Sail System (ACS3) CubeSat [72], and two simulating DLR’s PIXL-1 CubeSat [73]. We selected these two particular CubeSats because they use the CSP ecosystem (Sec. II-E) and are currently in-orbit and operational. We ensured these additional deployments were believable by configuring their satellite personalities and giving them public IP addresses in the regions belonging to the satellite mission origin, namely Germany and the United States. Table V summarizes all five HoneySat deployment details.
TABLE V: HoneySat Internet deployments’ details.
Deployment
Type
Satellite
Personality
HoneySat IP
Location
Duration
Cloud PIXL-1 Germany 6
Cloud PIXL-1 Germany 6
Cloud ACS3 United States 6
Cloud ACS3 United States 6
On-prem SUCHAI-2 University of Chile 12
TABLE VI: Exposed Telnet interactions received.
Satellite
Personality
HoneySat
IP
Attacker